OpenStudy (anonymous):

PLEASE HELP!!! Find and simplify the difference quotient f(x + h) − f(x)/h for the following function. f(x) = 8x^3

OpenStudy (anonymous):

Can you find \(f(x+h)\) first?

OpenStudy (anonymous):

umm is it 8x^3(x+h) ?

OpenStudy (anonymous):

No, what you wrote is \(f(x)\cdot (x+h)\)

OpenStudy (anonymous):

\[ f(x+h) = 8(x+h)^3 \]

OpenStudy (anonymous):

But you have to multiply it out.

OpenStudy (anonymous):

i forget how to do that when its a power to the third

OpenStudy (anonymous):

Can you do it to a power of 2?

OpenStudy (anonymous):

\[ 8(x+h)^3=8(x+h)^2(x+h) \]

OpenStudy (anonymous):

(8x^2+16xh+8h^2)(x+h)

OpenStudy (anonymous):

i think it is drivation formula and we should differentiate that and we'll have 24x^2

OpenStudy (debbieg):

This isn't a differentiation problem. He probably hasn't learned derivatives yet, you learn the difference quotient well before that. He needs to do the problem that was asked.

OpenStudy (anonymous):

well im a she btw.. haha

OpenStudy (debbieg):

Sorry, the name is a little bit hard to tell.... and that picture, well, I never can figure it out with this darn owls! ;) lol

OpenStudy (anonymous):

I knew it was a girl.

OpenStudy (anonymous):

haha its all good :p

OpenStudy (anonymous):

oh.. so sorry.. i don't know that

OpenStudy (anonymous):

so anyway whats the next step?

OpenStudy (anonymous):

@sara17 Even if you did the derivative, you'd lose the \(h\).

OpenStudy (anonymous):

The next step is to multiply out \[ (8x^2+16xh+8h^2)(x+h) \]

OpenStudy (anonymous):

i cant figure that part out

OpenStudy (anonymous):

yes but i think that in the end h will omit

OpenStudy (anonymous):

8x^3+16x^2h+8xh+8xh+16xh^2+8h^3

OpenStudy (anonymous):

??

OpenStudy (anonymous):

second and third power of h are insignificant and i think you should divide it by h

OpenStudy (anonymous):

i typed in your answer 24x^2 but it said it was wrong

OpenStudy (anonymous):

@wio am i right or wrong...

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

so ... which one is my answer.. the extended form?

OpenStudy (anonymous):

Expanded form.

OpenStudy (anonymous):

ugh it says thats wrong too

OpenStudy (anonymous):

That isn't the full solution. You need to subtract f(x)

OpenStudy (anonymous):

and divide by h.

OpenStudy (anonymous):

so i would subtract the expanded form - 8x^3

OpenStudy (anonymous):

: (8(x+h)^3-8x^3)/h (8x^3+8h^3+8x^2h+8h^2x+16h^2+16hx-8x^3) 8x^3+8h^3+24h^2x+24x^2h-8x^3/h 8h^2+24hx+24x^2 h is unsignificant so =24x^2

OpenStudy (anonymous):

its saying thats not the right answer. can someone just give me the final answer?

OpenStudy (anonymous):

the right answer is 8h^2+24hx+24x^2 but since h is come from the limitation concept so we can omit h and write it 24x^2

OpenStudy (anonymous):

You are wrong @sara17

OpenStudy (anonymous):

It never specifies what \(h\) is.

OpenStudy (mertsj):

\[8h^2+24xh+24x^2\]

OpenStudy (anonymous):

THANK YOU!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

so i'm not wrong that's exactly my answer

OpenStudy (debbieg):

@sara17, your answer was wrong because you keep saying that you can "omit h" which is not true at all, for a difference quotient. It isn't a limit, we haven't said anything about what h is, other than it is non-zero.